OpenStudy (ambermarie151):

what is the solution ? y=x+2 y=3x-2 A.(0,-2) B.(0,2) C.(4,2) D.(2,4)

OpenStudy (retireed):



OpenStudy (ambermarie151):

So it would be 2 divided by 4 which equals 2

OpenStudy (ambermarie151):

@retirEEd

OpenStudy (ijlal):

actually 4 divided by 2 equals 2 @ambermarie151 :P and yes kudos ! to @retirEEd for explaining :) Now simply plug in the value of x=2 in equation 1 or equation 2 i would plug it into equation y=x+2 y=2+2 y=4 so we found the point to be (2,4) Option D

OpenStudy (retireed):

@ijlal thanks for helping out, I was called to supper